OpenStudy (anonymous):

A tin can (right circular cylinder) with top and bottom is to have volume V . What dimensions (the radius of the bottom and the height) give the minimum total surface area?

OpenStudy (anonymous):

V=h(pi)r^2 , A=h2(pi)r + 2(pi)r^2

OpenStudy (anonymous):

\[A=2V/r+2πr^2\]

OpenStudy (anonymous):

Setting the derivative to zero give us \[r=\sqrt[3]{V/2 \pi}\]

OpenStudy (anonymous):

which equation should I use to find h